OpenStudy (anonymous):

what is the P(z<0.59)

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Z_table#Cumulative_table look up 0.59 to get the probability

OpenStudy (john_es):

Do you understand how to do it?

OpenStudy (anonymous):

just locate the value in the table? is that right?

OpenStudy (anonymous):

another question: computing for P(-2.0<=Z<=2.5) is the same as P(-2.0<Z<2.5)? Thanks!

OpenStudy (john_es):

Yes, true. :)

OpenStudy (anonymous):

the first equation in the ff up question is: \[-2.0 \le z \le 2.5\]

OpenStudy (john_es):

\[P(-2.0\leq z\leq 2.5)=P(z\leq2.5)-P(z\leq -2.0)\]

OpenStudy (john_es):

\[P(z\leq -2.0)=1-P(z\leq 2.0)\]

OpenStudy (anonymous):

i see. then add the two values?

OpenStudy (john_es):

Yes, this is the solution.

OpenStudy (anonymous):

is it the same if < is used instead of ≤?

OpenStudy (anonymous):

Thanks a lot for your help :)

OpenStudy (john_es):

In the normal distribution, "for practical effects" we can think is the same. It is not the same if we apply the rule to Binomial distributions (approximated by normal distributions).

OpenStudy (john_es):

In general, < and ≤ are not the same if we have discrete distributions (binomial, poisson, etc).

OpenStudy (anonymous):

Got it. Thanks!

OpenStudy (john_es):

You're welcome ;).

OpenStudy (anonymous):

Problem: suppose the total carb intake in 12 to 14-year old boys is normally distributed with mean 124g/1000cal and SD 20g/1000cal. What percentage of boys in this age range have carb intake above 140g/1000cal?

OpenStudy (john_es):

Taking X as the grams per calories, you should use, \[P(X\geq 0.14)\]Then you have to make the tipification, \[P\left(\frac{X-\mu}{\sigma}\geq\frac{0.14-0.124}{0.02}\right)\Rightarrow P\left(Z\geq\frac{0.14-0.124}{0.02}\right)\]where, \[\mu=\frac{124}{1000}\ g/cal\\ \sigma=\frac{140}{1000}\ g/cal\]

OpenStudy (john_es):

The result for the probability using the table and the properties I wrote before is, P=0.2119, and in percentage, 21.19%.

OpenStudy (anonymous):

Thank you! :)